# weaker gravitational waves

### v1

@@ 5 Numerical results: gravitational waves
Results of numerical simulations of GWs production by biased DWs with vacuum initial conditions are shown in Figs. 78, and 9.
Figure 7: GW spectra obtained with the $2048^{3}$ lattice for the bias parameter $\epsilon=0.025$ (top panel) and $\epsilon=0.05$ (bottom panel) starting with vacuum initial conditions and momentum cutoff $k_{cut}=1$.
Figure 8: GW spectra obtained at $2048^{3}$ lattice with vacuum initial conditions and momentum cutoff $k_{cut}=1$. The bias parameter equals $\epsilon=0.025$ (top left panel), $\epsilon=0.035$ (top right panel), $\epsilon=0.05$ (bottom left panel). Bottom right panel shows dependence of $\Omega_{gw,peak}$ on $\epsilon$. The spectra are produced with the fixed ratio $\kappa^{\prime}/\kappa=\pi/2$ (see description in the end of Sec. 3). This explains appearance of the artificial UV peak compared to Fig. 7.
@@
We restrict our analysis to relatively large values $\epsilon=0.025,0.035,0.05$, because otherwise it takes a long time for the DW network to collapse and hence for the GW spectrum to get stabilised. We have fixed other model parameters to be $\lambda=0.03$, $v=6\cdot 10^{16}~{}\mbox{GeV}$, and the number of relativistic degrees of freedom in the primordial plasma $g_{*}(T)=100$ (we keep it constant within the relevant timespan of GW production). The results can be rescaled to arbitrary values, as we demonstrate below. Figures 7 and 8 show spectra obtained with one simulation each on $2048^{3}$ lattice, while plot on Fig. 9 is produced by averaging over 10 simulations performed with different seeds for initial values of $\chi$ on $1024^{3}$ lattice.
@@
Figure 9: GW spectra obtained by averaging over 10 realizations assuming the bias parameter $\epsilon=0.05$ and vacuum initial conditions with momentum cutoff $k_{cut}=1$. Simulations have been performed with $1024^{3}$ lattice. The spectra have been produced with the fixed ratio $\kappa^{\prime}/\kappa=\pi/2$ (see description in the end of Sec. 3). This explains appearance of the artificial UV peak compared to Fig. 7. The quality of the fit to the IR slope is not good due to a lack of points.
@@
In this way, one can conservatively estimate the error bars related to the statistical nature of the initial field $\chi$ distribution. The spectrum of GWs in Fig. 7 is obtained with the procedure described in Sec. 3, cf. Ref. [27]. Namely, one varies the ratio $\kappa^{\prime}/\kappa$ according to Eq. (23), with parameters $\kappa^{\prime}$ and $\kappa$ controlling IR and UV physics, respectively. This allows us to “zoom in” the IR, near peak, and UV parts of the spectrum. Then one combines different parts to get the full spectrum. The procedure is designed to get rid of/mitigate the effect of non-physical factors. In particular, as it is clear from Fig. 8 obtained with fixed $\kappa^{\prime}/\kappa=\pi/2$, one observes the strong UV peak. The latter has a non-physical origin, as it has been proven in Ref. [27]: its location changes depending on the lattice resolution.
@@
which matches well the result obtained in the case of unbiased DWs [27]. The fractional energy density of GWs at maximum reconstructed from Figs. 7 and 8 is given by
@@
Note that the dependence on the expectation value $v$ and the constant $\lambda$, encoded in $H_{i}=\sqrt{\lambda}v$, follows from the relation $\rho_{gw}\propto\sigma^{2}_{wall}/M^{2}_{P}$. The factor $\sim H^{2}_{ann}$ in the denominator comes from the Friedmann equation, $\rho_{tot}=3H^{2}_{ann}M^{2}_{P}$, which relates the total energy density entering $\Omega_{gw}$ and the Hubble parameter. The peak energy density of GWs exceeds by about an order of magnitude the naive expectations, where $\tau_{ann}$ is the last instant of GW emission, cf. Eq. (58) of Ref. [27]. The deviation can be explained by the fact that GW production continues some time after the DW collapse $\tau_{ann}$. Namely, one has
@@
where $\tau_{gw}$ is the time of peak production of GW energy, cf. Ref. [19]. Indeed, one can see in Figs. 7 and 8, that the growth of GW energy density relative to radiation continues beyond the time $\tau_{ann}$. The possible physical reason is that the inhomogeneous particle distribution following the DW collapse continues to source GW waves. Such a source is terminated, once the spatial distribution of particles becomes homogeneous. If particles are relativistic, it takes about one Hubble time for the particle distribution to become homogeneous; since this time on the source of GWs is terminated, and their growth stops.
@@
Note that we have set $C(k_{cut})\approx 1$ in Eq. (27), because we assume the momentum cutoff $k_{cut}=1$ for vacuum initial conditions. The reference values in Eqs. (31) and (32) are chosen for convenience of comparison with the recent PTA findings [2, 3, 4, 5, 6, 7]. As it follows from Eq. (29), dependence of $\Omega_{gw,peak}$ on $\epsilon$ is considerably softer compared to the commonly assumed $\Omega_{gw,peak}\propto 1/\epsilon^{2}$. This is due to the earlier annihilation time revealed in our calculations as compared to the naively expected annihilation time. Consequently, one gets a much weaker GW signal for the same $\epsilon$, given that the latter is typically taken to be extremely small, as it is indicated in Eq. (32). This our conclusion is at odds with those of Refs. [17, 19, 20].
Now let us discuss the spectral shape of GWs. Spectral indices in Figs. 7 and 9 have been obtained using the following procedure. Assuming the power-law fit, so that $\ln\Omega_{gw}=\ln\alpha+\beta\ln k$ with constant $\alpha$ and $\beta$, we obtain values of $\alpha$ and $\beta$ and their errors using the linear least squares method. To estimate the spectral index in the IR part of the spectrum we take all the points from $k_{min}=2\pi/L$ to $k_{peak}$. In the UV part of the spectrum, we take all the points from $k=10\pi Ha$ to the momentum $k$, which is $1.5$ times smaller than the IR edge of the plateau. All the fits have been performed using function curve_fit function from Python library SciPy.
@@
i) The IR part of the spectrum, where one has $\Omega_{gw}\propto f^{3}$ as $f\rightarrow 0$ for causality considerations. Departures from this asymptotic behavior in Figs. 78, and 9 is due to the limited range of momenta available in simulations, i.e., $k\geq 2\pi/L$.
ii) The close-to-maximum UV part has the spectral shape approximately described by $\Omega_{gw}\propto f^{-1}$, which agrees with the results obtained recently in Ref. [17], but disagrees with those of Ref. [20]. Here one observes a rather strong deviation from the unbiased case characterized by the spectral shape $\Omega_{gw}\propto f^{-1.5}$ [27]. We attribute this feature to the richer substructure in the form of closed DWs in the case of biased DWs. Indeed, it has been observed in Refs. [27, 35] that the unbiased DW network is dominated by a single “infinite” wall, while the number of small closed walls is negligible. It is interesting to note that the spectral shape $\Omega_{gw}\propto f^{-1}$ follows from bubble collisions during the first order phase transitions [34].
iii) The far UV zone, where the GW spectrum exhibits a clear plateau. While a similar feature has been present in the case of unbiased DWs, here it is pronounced more sharply. The IR edge of the plateau is remarkably stable, independent of the simulation time and the bias parameter $\epsilon$. Given that one can relate the position of the IR edge of the plateau with the scale $k/a\sim 2\pi\sqrt{H/\delta_{wall}}$. This plateau is likely to have a physical origin. In particular, it becomes more pronounced once the role of the artificial UV peak is reduced. Furthermore, it has been shown in Ref. [27] that the plateau remains intact upon switching from the lattice $2048^{3}$ to $1024^{3}$ lattice; hence, it is unlikely to be an artefact of the limited lattice resolution at those small scales.
@@
where $x\equiv f/f_{peak}$, and $a,b,c$ are the fitting parameters. We have reconstructed the latter again using curve_fit function from SciPy Python library. Results are shown in Table 1, where we also compare the cases $\epsilon\neq 0$ with the unbiased case, $\epsilon=0$; fit to the average over 10 simulation spectrum is performed with $\chi^{2}_{\nu}=2.08$. We have taken GW spectrum produced by unbiased DWs from Fig. 1 in Ref. [27]; see the top panel there corresponding to vacuum initial conditions. We observe that introducing a bias leads to the decrease of power in the IR part of the spectrum and increase in the UV part.
@@ 6 Conclusions
The slight breaking of $Z_{2}$-symmetry is commonly introduced for solving the DW-domination problem. In the present work, we have numerically studied the impact of the potential bias $V_{bias}$, originating from the symmetry breaking, on the DW network evolution in a radiation-dominated Universe and the consequent production of GWs. It appears that the common way of determining the DW collapse time through the estimate $\sigma_{wall}H_{ann}\sim V_{bias}\sim\epsilon v^{3}$ may be misleading. A more accurate estimate reads (27), corresponding to the annihilation time (1). This difference, while looking relatively innocuous, can lead to dramatic results in the regime of very small $\epsilon$. In particular, due to an earlier network collapse compared to naive expectations, one expects a parametrically weaker GW signal for the same bias parameter $\epsilon$. Thus, in order to have an observable GW signal, one should assume a considerably smaller potential bias $\epsilon$ than previously thought. However, note that the $\epsilon$-dependence of the annihilation time has been inferred from a rather narrow range of values of $\epsilon$. In the future, owing to higher-resolution simulations and hence longer simulation times, one should be able to probe very small values of $\epsilon$ and establish the dependence $t_{ann}(\epsilon)$ on firm grounds.
@@
From the viewpoint of observations, the most interesting piece of information is contained in the GW power spectrum shape. While the spectral shape of GWs produced by the network of biased walls is similar to that in the case $\epsilon=0$, there are important differences most notably in the UV part, where GWs from biased DWs exhibit larger power compared to the unbiased case. In this regard, we confirm the results obtained recently in Ref. [17], which uses a different code, but disagree with those in Ref. [20], which exploits a modified CosmoLattice code. This tension is worth exploring in the future. In terms of the fitting formula (33), we have restricted values of the parameters $a,~{}b$ and $c$. Note that the plateau part of the spectrum originally observed in Refs. [19, 27] for unbiased DWs is also present in the biased case. Interestingly, in terms of conformal momentum, the position of IR edge of the plateau exhibits a remarkable stability, i.e., it remains constant as a function of time and potential bias. This may suggest that the plateau, if due to some physical mechanism rather than a simulation artefact, starts at $k/a\sim 2\pi\sqrt{H/\delta_{wall}}$. While there is currently no consensus on the nature of the plateau, our simulations in this paper and in Ref. [27] indeed suggest that the plateau has a physical origin.
@@
In the future, it will be interesting to analyze how biased DWs perform with respect to PTA data relatively to other GW sources. In particular, there is another type of DWs, i.e., melting DWs, capable of producing an observable GW signal [14, 15]. These melting DWs are described by a time-decreasing tension and they do not overclose the Universe, even in the case of an exact $Z_{2}$-symmetry. Spectral shapes of conventional constant tension (biased or unbiased) DWs and melting DWs are drastically different, especially in the close-to-maximum IR region [^3] At the same time in the close-to-maximum UV range, spectral shapes of GWs sourced by melting DWs and biased DWs are very similar. As in Sec. 5, we explain this similarity by noticing that this part of the spectrum is sensitive to closed DWs (analogous to vacuum bubbles during first order phase transitions). Indeed, it has been demonstrated in Ref. [36] that a large fraction of melting DW network is in the form of closed walls., where one has $\Omega_{gw}\propto f^{1.6}$ in the case of melting DWs [36]. While the current PTA data releases are more in favor of melting DWs [37], the error bars are presently too large to clearly discriminate between the two sources. We expect that future PTA data releases will provide further observational support for one of these exciting high energy physics scenarios.

### v2

@@ 5 Numerical results: gravitational waves
Results of numerical simulations of GWs production by biased DWs with vacuum initial conditions are shown in Figs. 78, and 9. We restrict our analysis to relatively large values $\epsilon=0.018,0.025,0.035,0.05$, because otherwise it takes a long time for the DW network to collapse and hence for the GW spectrum to get stabilised. We have fixed other model parameters to be $\lambda=0.03$, $v=6\cdot 10^{16}~{}\mbox{GeV}$, and the number of relativistic degrees of freedom in the primordial plasma $g_{*}(T)=100$ (we keep it constant within the relevant timespan of GW production). The results can be rescaled to arbitrary values, as we demonstrate below. Figures 7 and 9 show spectra obtained with one simulation each on $2048^{3}$ lattice, while the plot on Fig. 8 is produced by averaging over 10 simulations performed with different seeds for initial values of $\chi$ on $1024^{3}$ lattice.
Figure 7: GW spectra obtained with the $2048^{3}$ lattice for the bias parameter $\epsilon=0.025$ (top panel) and $\epsilon=0.05$ (bottom panel) starting with vacuum initial conditions and momentum cutoff $k_{cut}=1$. Shown also are power-law fits and the fit described by Eq. (33).
@@
Figure 8: GW spectra obtained by averaging over 10 realisations assuming the bias parameter $\epsilon=0.05$ and vacuum initial conditions with momentum cutoff $k_{cut}=1$. Simulations have been performed with $1024^{3}$ lattice. The spectra have been produced with the fixed ratio $\kappa^{\prime}/\kappa=\pi/2$ (see description in the end of Sec. 3). This explains appearance of the artificial UV peak compared to Fig. 7. Shown also are power-law fits and the fit described by Eq. (33). The quality of the fit to the IR slope is not good due to a lack of points.
Figure 9: GW spectra obtained at $2048^{3}$ lattice with vacuum initial conditions and momentum cutoff $k_{cut}=1$. The bias parameter equals $\epsilon=0.018$ (top left panel), $\epsilon=0.025$ (top right panel), $\epsilon=0.035$ (bottom left panel), $\epsilon=0.05$ (bottom right panel). The spectra are produced with the fixed ratio $\kappa^{\prime}/\kappa=\pi/2$ (see description in the end of Sec. 3). This explains appearance of the artificial UV peak compared to Fig. 7.
Figure 10: Demonstrated is dependence of $\Omega_{gw,peak}$ on the bias parameter $\epsilon$ extracted from Fig. 9. Vacuum initial conditions with momentum cutoff $k_{cut}=1$ have been assumed.
@@
In this way, one can conservatively estimate error bars related to the statistical nature of the initial field $\chi$ distribution. The spectrum of GWs in Fig. 7 is obtained with the procedure described in Sec. 3, cf. Ref. [28]. Namely, one varies the ratio $\kappa^{\prime}/\kappa$ according to Eq. (23), with parameters $\kappa^{\prime}$ and $\kappa$ controlling IR and UV physics, respectively. This allows us to “zoom in” the IR, near peak, and UV parts of the spectrum. Then one combines different parts to get the full spectrum. The procedure is designed to get rid of/mitigate the effect of non-physical factors. In particular, as it is clear from Figs. 8 and 9 obtained with fixed $\kappa^{\prime}/\kappa=\pi/2$, one observes the strong UV peak. The latter has a non-physical origin, as it has been proven in Ref. [28]: its location changes depending on the lattice resolution.
@@
which matches well the result obtained in the case of unbiased DWs [28]. The fractional energy density of GWs at maximum reconstructed from Figs. 7 and 9 is given by
@@
Note that the dependence on the expectation value $v$ and the constant $\lambda$, encoded in $H_{i}=\sqrt{\lambda}v$, follows from the relation $\rho_{gw}\propto\sigma^{2}_{wall}/M^{2}_{P}$. The factor $\sim H^{2}_{ann}$ in the denominator comes from the Friedmann equation, $\rho_{tot}=3H^{2}_{ann}M^{2}_{P}$, which relates the total energy density entering $\Omega_{gw}$ and the Hubble parameter. The peak energy density of GWs exceeds by about an order of magnitude the naive expectations, where $\tau_{ann}$ would be taken as the last instant of GW emission, cf. Eq. (58) of Ref. [28]. The deviation can be explained by the fact that GW production continues some time after the DW collapse $\tau_{ann}$. Namely, one has
@@
where $\tau_{gw}$ is the time of peak production of GW energy, cf. Refs. [16, 21]. Indeed, one can see in Figs. 7 and 9, that the growth of GW energy density relative to radiation continues beyond the time $\tau_{ann}$. The possible physical reason is that the inhomogeneous particle distribution following the DW collapse continues to source GW waves. Such a source is terminated, once the spatial distribution of particles becomes homogeneous. If particles are relativistic, it takes about one Hubble time for the particle distribution to become homogeneous; since this time on the source of GWs is terminated, and their growth stops.
@@
Note that we have set $C(k_{cut})\approx 1$ in Eq. (27), because we assume the momentum cutoff $k_{cut}=1$ for vacuum initial conditions. The reference values in Eqs. (31) and (32) are chosen for convenience of comparison with the recent PTA findings [2, 3, 4, 5, 6, 7]. As it follows from Eq. (29), see also Fig. 10, dependence of $\Omega_{gw,peak}$ on $\epsilon$ is considerably softer compared to the commonly assumed $\Omega_{gw,peak}\propto 1/\epsilon^{2}$. This is due to the earlier annihilation time revealed in our calculations as compared to the naively expected annihilation time. Consequently, one gets a much weaker GW signal for the same $\epsilon$, given that the latter is typically taken to be extremely small, as it is indicated in Eq. (32). This conclusion of our analysis is at odds with those of Refs. [17, 21, 18].
Now let us discuss the spectral shape of GWs. Spectral indices in Figs. 7 and 8 have been obtained using the following procedure. Assuming the power-law fit, so that $\ln\Omega_{gw}=\ln\alpha+\beta\ln k$ with constant $\alpha$ and $\beta$, we obtain values of $\alpha$ and $\beta$ and their errors using the linear least squares method. To estimate the spectral index in the IR part of the spectrum we take all the points from $k_{min}=2\pi/L$ to $k_{peak}$. In the UV part of the spectrum, we take all the points from $k_{peak}$ to the momentum $k$, which is $1.5$ times smaller than the IR edge of the plateau. All the fits have been performed using function curve_fit function from Python library SciPy. The choice of conformal times to fit GW spectra is dictated by the following considerations: the spectra should be stabilised at those times, and oscillatory features imposed on the spectra should not be prominent. Note that oscillations must be averaged out anyway, when calculating the GW energy density. Such an averaging can be trivially implemented, if the GW period is small relative to the Hubble radius. Due to a limited timespan of our simulations, however, certain modes of interest have wavelengths of the order of the horizon radius by the end of simulations, and it may be difficult to get rid of oscillations in the IR part of the spectrum. In this situation, one mitigates oscillations simply by picking the spectrum at the time, when oscillatory effects are accidentally small.
@@
i) The IR part of the spectrum, where one has $\Omega_{gw}\propto f^{3}$ as $f\rightarrow 0$ for causality considerations. Departures from this asymptotic behavior in Figs. 78, and 9 are due to the limited range of momenta available in simulations, i.e., $k\geq 2\pi/L$.
ii) The close-to-maximum UV part has the spectral shape approximately described by $\Omega_{gw}\propto f^{-1}$, which agrees with the results obtained recently in Ref. [17], but disagrees with those of Ref. [18]. Here one observes a rather strong deviation from the unbiased case characterised by the spectral shape $\Omega_{gw}\propto f^{-1.5}$ [28]. We attribute this feature to the richer substructure in the form of closed DWs in the case of biased DWs. Indeed, it has been observed in Refs. [28, 36] that the unbiased DW network is dominated by a single “infinite” wall, while the number of small closed walls is negligible. It is interesting to note that the spectral shape $\Omega_{gw}\propto f^{-1}$ follows from bubble collisions during the first order phase transitions [35].
iii) The far UV zone, where the GW spectrum exhibits a clear plateau. While a similar feature has been present in the case of unbiased DWs, here it is pronounced more sharply. The IR edge of the plateau is remarkably stable: it is independent of the simulation time and the bias parameter $\epsilon$. Given that one can relate the position of the IR edge of the plateau with the scale $k/a\sim 4\pi\sqrt{H/\delta_{wall}}$, this plateau is likely to have a physical origin. In particular, it becomes more pronounced once the role of the artificial UV peak is reduced. Furthermore, it has been shown in Ref. [28] that the plateau remains intact upon switching from the lattice $2048^{3}$ to $1024^{3}$ lattice; hence, it is unlikely to be an artefact of the limited lattice resolution at those small scales.
@@
where $x\equiv f/f_{peak}$, and $a,b,c$ are the fitting parameters. We have reconstructed the latter again using curve_fit function from SciPy Python library. The fitting procedure has been applied to GW spectra with oscillatory features suppressed, as in the case of power-law fits discussed above. Results are shown in Table 1, where we also compare the cases $\epsilon\neq 0$ with the unbiased case, $\epsilon=0$; fit to the average over 10 simulation spectrum is performed with $\chi^{2}_{\nu}=1.63$. We have taken GW spectrum produced by unbiased DWs from Fig. 1 in Ref. [28]; see the top panel there corresponding to vacuum initial conditions. We observe that introducing a bias leads to the decrease of power in the IR part of the spectrum and increase in the UV part.
@@ 6 Conclusions
The slight breaking of $Z_{2}$-symmetry is commonly introduced for solving the DW-domination problem. In the present work, we have numerically studied the impact of the potential bias $V_{bias}$, originating from the symmetry breaking, on the DW network evolution in a radiation-dominated Universe and the consequent production of GWs. It appears that the common way of determining the DW collapse time through the estimate $\sigma_{wall}H_{ann}\sim V_{bias}\sim\epsilon v^{3}$ may be misleading. A more accurate estimate reads (27), corresponding to the annihilation time (1). This difference, while looking relatively innocuous, can lead to dramatic results in the regime of very small $\epsilon$. In particular, due to an earlier network collapse compared to naive expectations, one expects a parametrically weaker GW signal for the same bias parameter $\epsilon$. Thus, in order to have an observable GW signal, one should assume a considerably smaller potential bias $V_{bias}\propto\epsilon$ than previously thought. However, note that the $\epsilon$-dependence of the annihilation time has been inferred from a rather narrow range of values of $\epsilon$. In the future, owing to higher-resolution simulations and hence longer simulation times, one should be able to probe very small values of $\epsilon$ and establish the dependence $t_{ann}(\epsilon)$ on firm grounds.
@@
From the viewpoint of observations, the most interesting piece of information is contained in the GW power spectrum shape. While the spectral shape of GWs produced by the network of biased walls is similar to that in the case $\epsilon=0$, there are important differences most notably in the UV part, where GWs from biased DWs exhibit larger power compared to the unbiased case. In this regard, we confirm the results obtained recently in Ref. [17], which uses a different code, but disagree with those in Ref. [18], which exploits a modified CosmoLattice code. This tension is worth exploring in the future. In terms of the fitting formula (33), we have restricted values of the parameters $a,~{}b$ and $c$. Note that the plateau part of the spectrum observed in Refs. [27, 28] for unbiased DWs is also present in the biased case, cf. Ref. [16]. Interestingly, in terms of conformal momentum, the position of IR edge of the plateau exhibits a remarkable stability, i.e., it remains constant as a function of time and potential bias. This may suggest that the plateau, if due to some physical mechanism rather than a simulation artefact, starts at $k/a\sim 4\pi\sqrt{H/\delta_{wall}}$. While there is currently no consensus on the nature of the plateau, our simulations in this paper and in Ref. [28] indeed suggest that the plateau has a physical origin.
@@
In the future, it will be interesting to analyse how biased DWs perform with respect to PTA data relatively to other GW sources. In particular, there is another type of DWs, i.e., melting DWs, capable of producing an observable GW signal [14, 15]. These melting DWs are described by a time-decreasing tension and they do not overclose the Universe, even in the case of an exact $Z_{2}$-symmetry. Spectral shapes of conventional constant tension (biased or unbiased) DWs and melting DWs are drastically different, especially in the close-to-maximum IR region [^3] At the same time in the close-to-maximum UV range, spectral shapes of GWs sourced by melting DWs and biased DWs are very similar. As in Sec. 5, we explain this similarity by noticing that this part of the spectrum is sensitive to closed DWs (analogous to vacuum bubbles during first order phase transitions). Indeed, it has been demonstrated in Ref. [37] that a large fraction of melting DW network is in the form of closed walls., where one has $\Omega_{gw}\propto f^{1.6}$ in the case of melting DWs [37]. While the current PTA data releases are more in favor of melting DWs [38], the error bars are presently too large to clearly discriminate between the two sources. We hope that future PTA data releases will provide further observational support for one of these high energy physics scenarios.